Two cars glide without friction along a level air track.

Before collision:

Car 1 has mass 294 g and velocity 2.75 m/s to the right.

Car 2 has mass 264 g and is at rest.

After a perfectly elastic collision, find the final velocity of Car 2 where right is positive.
Answer #1

Take right direction as positive and left as negative.

using conservation of momentum

m1v1i + m2v2i = m1v1f + m2v2f

as car 2 was initially at rest, v2i = 0

the equations which will give us final velocity of car 1 is

v1f = ( m1 - m2) * v1i / m1 + m2

v1f = (0.294 - 0.264 / 0.294 + 0.264 ) * 2.75

v1f = 0.1478 m/s

similarly,

v2f = (2m1 / m1 + m2 )v1i

v2f = 2.897 m/s